# A Nonabelian Hodge Correspondence for Principal Bundles in Positive Characteristic

### v1

@@ 2.3. Stability Conditions
A flat $G$-bundle $(V,\nabla)$ is $R$-semistable (resp. $R$-stable), if
  • for any proper parabolic subgroup $P\subseteq G$;
  • for any antidominant character $\chi:P\rightarrow\mathbb{G}_{m}$ acting trivially on the center;
  • for any $\nabla$-compatible reduction of structure group $\varsigma:X\rightarrow V/P$,
we have
[…]
Let ${\rm MIC}^{(s)s}_{p-1}(X/k,G)$ be the category of $R$-(semi)stable nilpotent flat $G$-bundles of exponent $\leq p-1$.
@@ 3. Nilpotent $G$-Higgs Bundles and Nilpotent Flat $G$-Bundles
The goal of this section is proving the equivalence of categories under the $W_{2}(k)$-liftable condition:
@@
To prove the theorem, we construct inverse Cartier functor in §3.1
@@
via an approach called exponential twisting introduced in [LSZ15]. These two functors are inverse to each other, and then induces the equivalence of categories. Furthermore, the equivalence holds under stability conditions. In other words, categories ${\rm HIG}^{(s)s}_{p-1}(X^{\prime}/k,G)$ and ${\rm MIC}^{(s)s}_{p-1}(X/k,G)$ are equivalent (Theorem 3.6). The key point to generalize Lan–Sheng–Zuo’s approach to principal bundles is to give an appropriate definition of the exponentiation in the case of principal bundles, which has been done by a series work (see [Ser94, Sei00, McN02, Sob15] for instance) and reviewed in §2. Before we move to the Cartier and inverse Cartier transform, we first review two classical results: classical Cartier descent and Deligne-Illusie’s lemma, which will be used frequently in this section.
@@
Given $(E,0)\in{\rm HIG}_{0}(X^{\prime}/k)$, Katz constructed a canonical connection $\nabla_{can}$ on $F^{*}_{X/k}E$, which induces a functor ${\rm HIG}_{0}(X^{\prime}/k)\rightarrow{\rm MIC}_{0}(X/k)$, and prove that the categories are equivalent. Chen–Zhu extended this result to principal bundles:
@@
Let $\mathcal{G}^{\prime}$ be a smooth affine group scheme over $X^{\prime}$ and define $\mathcal{G}:=F^{*}_{1}\mathcal{G}^{\prime}$. Then we have an equivalence of categories between the category of $\mathcal{G}^{\prime}$-torsors on $X^{\prime}$ and the category of $\mathcal{G}$-local systems with trivial $p$-curvature. As a special case, when $\mathcal{G}^{\prime}=X^{\prime}\times G$, then the categories ${\rm HIG}_{0}(X^{\prime}/k,G)$ and ${\rm MIC}_{0}(X/k,G)$ are equivalent.
@@ A Lemma by Deligne-Illusie
We review a lemma by Deligne–Illusie and we refer the reader to [DI87] and [LSZ15] for more details. Let $W_{2}(k)$ be the ring of Witt vectors of length two. We fix a $W_{2}(k)$-lifting $X_{2}$ of $X$, and choose an open affine cover $\{U_{2,i}\}_{i\in I}$ of $X_{2}$, where $I$ is the index set. Let $F_{2,i}:U_{2,i}\rightarrow U^{\prime}_{2,i}$ be a Frobenius lifting, which ${\rm mod}\,p$ is the relative Frobenius morphism $F_{1,i}:U_{i}\rightarrow U^{\prime}_{i}$, and we use the notation $F_{1}:=F_{1,i}$ for convenience. We have a natural morphism
@@ 3.1. Inverse Cartier Transform
In this section, we will show that given a nilpotent $G$-Higgs bundle $(E,\theta)$, it can be associated with a nilpotent flat $G$-bundle. Furthermore, this construction induces a functor
We fix some notations first. Let $\mathcal{U}=\{U_{i}\}_{i\in I}$ be a covering of open affine subsets of $X$ given in Lemma 3.4. Given a $G$-Higgs bundle $(E,\theta)$ on $X^{\prime}$, denote by $E_{i}:=E|_{U_{i}}$ the restriction of $E$ to the open set $U_{i}$. We can define the above notations on the intersections $U_{ij}$ similarly. For convenience, we use the same notation $F_{1}$ for relative Frobenius maps of $X$, $U_{i}$ and etc. for simplicity.
Let $(E,\theta)$ be a $G$-Higgs bundle in ${\rm HIG}_{p-1}(X^{\prime}/k,G)$. The element $h_{ij}(F_{1}^{*}\theta)$ is a section of $(F_{1}^{*}E_{ij})(\mathfrak{g})$. Since $\theta$ is nilpotent of exponent $\leq p-1$, the element
$G_{ij}:=\overline{\exp}(h_{ij}(F_{1}^{*}\theta))$
is well-defined by Theorem 2.5 and induces an automorphism of $F^{*}_{1}E_{ij}$. Since $\theta$ is integrable, we have
@@
By gluing $\{V_{i}:=F^{*}_{1}E_{i}\}_{i\in I}$ via $G_{ij}$, we obtain a $G$-bundle $V$ on $X$.
Next, we construct a connection on $V$. Denote by $\nabla_{can,i}$ the integrable connection on $F^{*}_{1}E_{i}$ given by Corollary 3.3, and we define a new connection
@@
$dG_{ij}\cdot G_{ij}^{-1}+{\rm Ad}(G_{ij})(\zeta_{j}(F^{*}_{1}\theta))=\zeta_{i}(F^{*}_{1}\theta).$
The first term $G_{ij}^{-1}\cdot dG_{ij}$ is $dh_{ij}(F_{1}^{*}\theta)$. For the second term ${\rm Ad}(G_{ij})(\zeta_{j}(F^{*}_{1}\theta))$, we have
${\rm Ad}(G_{ij})(\zeta_{j}(F^{*}_{1}\theta))={\rm Ad}(\overline{\exp}(h_{ij}(F_{1}^{*}\theta)))(\zeta_{j}(F^{*}_{1}\theta))=\zeta_{j}(F^{*}_{1}\theta)$
by the second property of the exponential map (Definition 2.4). Summing up, we have
$dG_{ij}\cdot G_{ij}^{-1}+{\rm Ad}(G_{ij})(\zeta_{j}(F^{*}_{1}\theta))=dh_{ij}(F_{1}^{*}\theta)+\zeta_{j}(F^{*}_{1}\theta).$
Applying Lemma 3.4 again, we get the desired equality and obtain a $G$-connection $\nabla$ on $V$.
@@
In conclusion, we obtain a functor
@@ 3.2. Cartier Transform
Given a pair $(V,\nabla)\in{\rm MIC}_{p-1}(X/k,G)$, denote by $\psi$ the $p$-curvature of $\nabla$. Note that $\psi$ is nilpotent of exponent $\leq p$ by definition. We use the same notation for coverings as in §3.1, and let $V_{i}:=V|_{U_{i}}$, $\psi_{i}:=\psi|_{U_{i}}$. Then, we define a new connection $\nabla^{\prime}_{i}$ on $V_{i}$ as
@@
Following the same argument as [LSZ15, §2.3, Step 3], the $p$-curvature of $\nabla^{\prime}_{i}$ is trivial. Moreover, since the $p$-curvature $\psi_{i}$ is horizontal with respect to $\nabla_{i}$, it is also horizontal with respect to $\nabla^{\prime}_{i}$. Now we want to glue $\{(V_{i},\nabla^{\prime}_{i},\psi_{i})\}_{i\in I}$ to obtain a new flat $G$-bundle $(V^{\prime},\nabla^{\prime})$, of which the $p$-curvature is trivial, together with a $F$-Higgs field $\psi$ horizontal with respect to $\nabla^{\prime}$. We define
@@
which is an automorphism of $V_{ij}$. With the same argument as formula (3.1), we have

### v2

@@ 2.2. Nilpotency
A $G$-Higgs bundle $(E,\theta)$ (resp. flat $G$-bundle $(V,\nabla)$) is said to be nilpotent of exponent $\leq n$ if there exists a covering of $X$ by open affine subsets $U_{i}$ such that for each $U_{i}$, the set
$\{\theta|_{U_{i}}(\partial)\ |\ \partial\ \text{is a section of }T_{U_{i}}\}\quad(resp.\text{ }\{\psi|_{U_{i}}(\partial)\ |\ \partial\ \text{is a section of }F^{*}_{U^{\prime}_{i}/k}T_{U^{\prime}_{i}}\})$
lies in $\mathfrak{u}(\mathcal{O}_{U_{i}})$, where $\mathfrak{u}$ is the Lie algebra of a unipotent radical of a parabolic subgroup $P$ of $G$ such that the nilpotency class of the unipotent radical is smaller than $n+1$.
@@ 3. Nilpotent $G$-Higgs Bundles and Nilpotent Flat $G$-Bundles
and suppose that the characteristic $p$ is good.
Let $G$ be a connected split reductive linear algebraic group. We introduce the following categories:
  • ${\rm HIG}_{0}(X/k,G)$: the category of $G$-bundles on $X$;
  • ${\rm MIC}_{0}(X/k,G)$: the category of flat $G$-bundles on $X$ with trivial $p$-curvature;
  • ${\rm HIG}_{p-1}(X/k,G)$: the category of nilpotent Higgs bundles on $X$ of exponent $\leq p-1$;
  • ${\rm MIC}_{p-1}(X/k,G)$: the category of nilpotent flat $G$-bundles on $X$ of exponent $\leq p-1$.
Let $\mathcal{C}$ be a category given above. We define $\mathcal{C}^{s(s)}$ to be the full subcategory of $\mathcal{C}$ of $R$-(semi)stable objects. For instance, ${\rm HIG}^{s(s)}_{p-1}(X/k,G)$ is the category of $R$-(semi)stable nilpotent $G$-Higgs bundles on $X$ of exponent $\leq p-1$.
The goal of this section is proving an equivalence of categories:
@@
Suppose that $X$ is $W_{2}(k)$-liftable and the characteristic $p$ is good. The categories ${\rm HIG}_{p-1}(X^{\prime}/k,G)$ and ${\rm MIC}_{p-1}(X/k,G)$ are equivalent.
The $W_{2}(k)$-liftable condition and the goodness of characteristics will be explained later. To prove the theorem, we construct inverse Cartier transform (as a functor) in §3.1
@@
via an approach called exponential twisting introduced in [LSZ15]. These two functors are inverse to each other, and then induces an equivalence of categories. Furthermore, the equivalence holds under stability conditions (Theorem 3.8).
Before we move to the construction of inverse Cartier and Cartier transform, we first review three classical results: classical Cartier descent, Deligne-Illusie’s lemma and exponentiation, which will be used in the proof.
@@
Given $(E,0)\in{\rm HIG}_{0}(X^{\prime}/k)$, Katz constructed a canonical connection $\nabla_{can}$ on $F^{*}_{X/k}E$ and prove that the induced functor ${\rm HIG}_{0}(X^{\prime}/k)\rightarrow{\rm MIC}_{0}(X/k)$ gives an equivalence of categories. Chen–Zhu extended this result to principal bundles:
@@
Let $\mathcal{G}^{\prime}$ be a smooth affine group scheme over $X^{\prime}$ and define $\mathcal{G}:=F^{*}_{1}\mathcal{G}^{\prime}$. Then we have an equivalence between the category of $\mathcal{G}^{\prime}$-torsors on $X^{\prime}$ and the category of $\mathcal{G}$-local systems with trivial $p$-curvature. As a special case, when $\mathcal{G}^{\prime}=X^{\prime}\times G$, the categories ${\rm HIG}_{0}(X^{\prime}/k,G)$ and ${\rm MIC}_{0}(X/k,G)$ are equivalent.
@@
Let $W_{2}(k)$ be the length two Witt vector ring. We fix a $W_{2}(k)$-lifting $X_{2}$ of $X$ and choose an affine open cover $\{U_{2,i}\}_{i\in I}$ of $X_{2}$, where $I$ is the index set. Let $F_{2,i}:U_{2,i}\rightarrow U^{\prime}_{2,i}$ be a Frobenius lifting, which ${\rm mod}\,p$ is the relative Frobenius morphism $F_{1,i}:U_{i}\rightarrow U^{\prime}_{i}$, and we use the notation $F_{1}:=F_{1,i}$ for simplicity. We have a natural morphism
@@
Exponentiation
Sobaje gave the definition of exponential maps for commuting nilpotent varieties [Sob15, Definition 2.3] and proved the uniqueness of the exponential maps when the characteristic $p$ is good. We refer the reader to [Her13, §2.9] for the definition of good characteristics.
Theorem 3.5 (Theorem 3.4 in [Sob15]).
Suppose that the characteristic $p$ is good. Let $\mathcal{N}_{1}(\mathfrak{g})$ be the $[p]$-nilpotent variety of $\mathfrak{g}$ and let $\mathcal{U}_{1}(G)$ be the $p$-unipotent variety of $G$. Then there is a unique $G$-equivariant bijection
Remark 3.6.
When the nilpotency class of $U$ is less than $p$, the $P$-equivariant isomorphism $\mathfrak{u}\xrightarrow{\cong}U$ is first studied by Serre [Ser94, §4] and then proved by Seitz [Sei00, §5], where Seitz proved the existence and uniqueness of the exponential map $\mathfrak{u}\rightarrow U$ via an argument of mod $p$ reduction. Since this approach is used in this section, we briefly review it as follows. We fix a maximal torus $T$ of $G$ which is contained in $P$, and let $\mathcal{R}^{+}$ be the set of positive roots. Let $\Gamma\subset\mathcal{R}^{+}$ be a closed subset of positive roots and denote by $U_{\Gamma}$ the corresponding unipotent subgroup with Lie algebra $\mathfrak{u}_{\Gamma}$. Let $i:{\rm Spec}\,k\rightarrow{\rm Spec}\,\mathbb{Z}_{(p)}$ be the composite ${\rm Spec}\,k\rightarrow{\rm Spec}\,\mathbb{F}_{p}\rightarrow{\rm Spec}\,\mathbb{Z}_{(p)}$. Then the exponential isomorphism $\overline{\rm exp}:\mathfrak{u}_{\Gamma}\rightarrow U_{\Gamma}$ in characteristic $p$ comes from the base change of the exponential isomorphism
$\widetilde{\rm exp}:\widetilde{\mathfrak{u}}_{\Gamma}\rightarrow\widetilde{U}_{\Gamma}$
in characteristic 00, where $\widetilde{U}_{\Gamma}$ (resp. $\widetilde{\mathfrak{u}}_{\Gamma}$) is the unipotent subgroup (resp. nilpotent Lie algebra) over $\mathbb{Z}_{(p)}$ such that $U_{\Gamma}=\widetilde{U}_{\Gamma}\times_{{\rm Spec}\,\mathbb{Z}_{(p)}}{\rm Spec}\,k$. We refer the reader to [Sei00, §5] for more details. Moreover, we can identifies $U_{\Gamma}$ with the affine space with coordinates $(x_{\alpha})_{\alpha\in\Gamma}$, then the exponential isomorphism can be expressed by the formula
$\overline{\rm exp}(D)=\bigg{(}\sum_{0\leq m\leq p-1}\dfrac{D^{m}}{m!}(x_{\alpha})\ \text{at}\ e\bigg{)}_{\alpha\in\Gamma},$
and we refer the reader to [BDP17, §6] for more details.
@@ 3.1. Inverse Cartier transform
In this subsection, we will show that given a nilpotent $G$-Higgs bundle of exponent $\leq p-1$, it can be associated with a nilpotent flat $G$-bundle of exponent $\leq p-1$. Furthermore, this construction induces a functor
We fix some notations first. Let $\mathcal{U}=\{U_{i}\}_{i\in I}$ be a covering of open affine subsets of $X$ given in Lemma 3.4. Given a $G$-Higgs bundle $(E,\theta)$ on $X^{\prime}$, denote by $(E_{i},\theta_{i})$ the restriction of $(E,\theta)$ to $U^{\prime}_{i}$. We can define the above notations on the intersections $U_{ij}$ similarly. For convenience, we use the same notation $F_{1}$ for relative Frobenius morphisms of $X$, $U_{i}$ and etc. for simplicity.
Let $(E,\theta)$ be a $G$-Higgs bundle in ${\rm HIG}_{p-1}(X^{\prime}/k,G)$. The element $h_{ij}(F_{1}^{*}\theta):=h_{ij}(F_{1}^{*}\theta_{ij})$ is a section of $(F_{1}^{*}E_{ij})(\mathfrak{g})$. Since $\theta$ is nilpotent of exponent $\leq p-1$, the set $\{\theta_{ij}(\partial)\ |\ \partial\ \text{is a section of }T_{U^{\prime}_{ij}}\}$ lies in a Lie algebra $\mathfrak{u}(\mathcal{O}_{U^{\prime}_{ij}})$, which is the Lie algebra of a unipotent subgroup $U$. Then by Theorem 3.5, there exists an isomorphism $\overline{\rm exp}:\mathfrak{u}\xrightarrow{\cong}U$ given by the exponentiation. Therefore, the element
$G_{ij}:=\overline{\exp}(h_{ij}(F_{1}^{*}\theta_{ij}))$
is well-defined and induces an automorphism of $F^{*}_{1}E_{ij}$. If there is no ambiguity, we use the notation
$h_{ij}(F_{1}^{*}\theta):=h_{ij}(F_{1}^{*}\theta_{ij})$
for simplicity. Since $\theta$ is integrable, we have
@@
Gluing $\{V_{i}:=F^{*}_{1}E_{i}\}_{i\in I}$ via $G_{ij}$, we obtain a $G$-bundle $V$ on $X$.
Next, we construct a connection on $V$. Denote by $\nabla_{can,i}$ the canonical connection on $F^{*}_{1}E_{i}$ given by Corollary 3.3, and we define a new connection
@@
$G_{ij}^{-1}\cdot dG_{ij}+{\rm Ad}(G_{ij})(\zeta_{j}(F^{*}_{1}\theta))=\zeta_{i}(F^{*}_{1}\theta).$
We claim
$G^{-1}_{ij}\cdot dG_{ij}=dh_{ij}(F^{*}_{1}\theta),\quad{\rm Ad}(G_{ij})(\zeta_{j}(F^{*}_{1}\theta))=\zeta_{j}(F^{*}_{1}\theta),$
and then, the left-hand side of equation is
($\ast$) $G_{ij}^{-1}\cdot dG_{ij}+{\rm Ad}(G_{ij})(\zeta_{j}(F^{*}_{1}\theta))=dh_{ij}(F^{*}_{1}\theta)+\zeta_{j}(F^{*}_{1}\theta).$
Therefore, the desired equality follows directly from Lemma 3.4 and we obtain a $G$-connection $\nabla$ on $V$. The claim can be proved by a mod $p$ reduction argument and we only give the proof of the equality
$G^{-1}_{ij}\cdot dG_{ij}=dh_{ij}(F^{*}_{1}\theta)$
as an example. Let $\rho:G\hookrightarrow{\rm GL}(V)$ be a faithful finite dimensional representation given by the regular representation (see [Hum75, §8.6] or [Mil17, Theorem 4.9] for instance). Let $m$ be an integer such that $p^{m}>n$ and we take a $W_{m}(k)$-lifting $\widetilde{U}_{ij}$ of $U_{ij}$. Note that such a lifting always exists because $U_{ij}$ is affine. Moreover, there also exists a $W_{m}(k)$-lifting $\widetilde{\rho}:\widetilde{G}\hookrightarrow{\rm GL}(\widetilde{V})$ by the construction of $\rho$. Then we have the following equality for the corresponding Lie algebras
$p\widetilde{\mathfrak{g}}=(p\mathfrak{gl}(\widetilde{V}))\cap\widetilde{\mathfrak{g}}$
because the induced morphism $d\widetilde{\rho}:\widetilde{\mathfrak{g}}\rightarrow\mathfrak{gl}(\widetilde{V})$ of Lie algebras is injective. Let $\theta_{ij}$ be the restriction of the Higgs field $\theta$ to $U_{ij}$. We choose a $W_{m}(k)$-lifting $\widetilde{\theta}_{ij}$ of $\theta_{ij}$. The equality ($\ast$3.1) in characteristic $p$ is equivalent to the following equality:
$\widetilde{G}_{ij}^{-1}\cdot d\widetilde{G}_{ij}\equiv d\widetilde{h}_{ij}(\widetilde{F}^{*}_{1}\widetilde{\theta}_{ij})\text{ mod }\left(p(\widetilde{\mathfrak{g}}\otimes_{{\rm Spec}\,W_{m}(k)}\Omega_{\widetilde{U}_{ij}})(\mathcal{O}_{\widetilde{U}_{ij}})\right),$
where $\widetilde{h}_{ij}$ (resp. $\widetilde{F}^{*}_{1})$ is the lifting of $h_{ij}$ (resp. $F^{*}_{1}$) and $\widetilde{G}_{ij}=\widetilde{\rm exp}(\widetilde{h}_{ij}(\widetilde{F}^{*}_{1}\widetilde{\theta}_{ij}))$. By the facts that the exponential map commutes with $d\widetilde{\rho}$ and the equality $p\widetilde{\mathfrak{g}}=(p\mathfrak{gl}(\widetilde{V}))\cap\widetilde{\mathfrak{g}}$, it is equivalent to show the following equality in the case of ${\rm GL}(V)$
${\rm exp}(A)^{-1}\cdot d{\rm exp}(A)\equiv dA\text{ mod }\left(p(\mathfrak{gl}(\widetilde{V})\otimes_{{\rm Spec}\,W_{m}(k)}\Omega_{\widetilde{U}_{ij}})(\mathcal{O}_{\widetilde{U}_{ij}})\right),$
where $A=d\widetilde{\rho}(\widetilde{h}_{ij}(\widetilde{F}^{*}_{1}\widetilde{\theta}_{ij}))\in\mathfrak{gl}(\widetilde{V})(\mathcal{O}_{\widetilde{U}_{ij}})$ and ${\rm exp}(A)=\sum\limits_{0\leq m\leq n-1}\dfrac{A^{m}}{m!}$. Then we have
$dA^{m}\equiv mA^{m-1}dA\text{ mod }\left(p(\mathfrak{gl}(\widetilde{V})\otimes_{{\rm Spec}\,W_{m}(k)}\Omega_{\widetilde{U}_{ij}})(\mathcal{O}_{\widetilde{U}_{ij}})\right).$
Therefore,
${\rm exp}(A)^{-1}\cdot d{\rm exp}(A)$ $\equiv{\rm exp}(-A)(\sum\limits_{0\leq m\leq n-2}\dfrac{A^{m}}{m!}dA)\equiv{\rm exp}(-A){\rm exp}(A)dA$
$\equiv dA\text{ mod }\left(p(\mathfrak{gl}(\widetilde{V})\otimes_{{\rm Spec}\,W_{m}(k)}\Omega_{\widetilde{U}_{ij}})(\mathcal{O}_{\widetilde{U}_{ij}})\right).$
This finishes the proof of the equality ($\ast$3.1). We remark here that we can not show this equality directly by choose a faithful representation of $G$ in characteristic $p$ since the exponential map may not commute with morphisms of algebraic groups in characteristic $p$.
@@
In conclusion, the above construction gives a functor
@@ 3.2. Cartier transform
Given a pair $(V,\nabla)\in{\rm MIC}_{p-1}(X/k,G)$, denote by $\psi$ the $p$-curvature of $\nabla$. Note that $\psi$ is nilpotent of exponent $\leq p-1$ by definition. We use the same notation for coverings as in §3.1, and let $V_{i}:=V|_{U_{i}}$, $\psi_{i}:=\psi|_{U_{i}}$. Then, we define a new connection $\nabla^{\prime}_{i}$ on $V_{i}$ as
@@
We regard the new connection as $\nabla^{\prime}_{i}:\mathcal{O}_{V_{i}}\rightarrow\mathcal{O}_{V_{i}}\otimes\Omega_{U_{i}}$ (see Remark 2.3), and then the $p$-curvature of $\nabla^{\prime}_{i}$ is trivial by following the same argument as in [LSZ15, §2.3, Step 3, page 866-869]. For completeness of the proof, we also include it here. Consider the algebra
$F_{1}^{*}(S^{<p}T_{U^{\prime}_{i}}):=\bigoplus_{0\leq m<p}F_{1}^{*}(S^{m}T_{U^{\prime}_{i}})=F_{1}^{*}(S\dot{\ }T_{U^{\prime}_{i}})/I,$
where $I$ is the ideal $F_{1}^{*}(S^{\geq p}T_{U^{\prime}_{i}}):=\bigoplus\limits_{m\geq p}F_{1}^{*}(S^{m}T_{U^{\prime}_{i}})$. There exists a connection $\nabla_{T}$ on $F_{1}^{*}(S^{<p}T_{U^{\prime}_{i}})$ with $p$-curvature $\psi_{\nabla_{T}}(\partial)(\tau)=\partial\tau$ for $\partial\in F_{1}^{*}T_{U^{\prime}_{i}},\ \tau\in F_{1}^{*}(S^{<p}T_{U^{\prime}_{i}})$ by [LSZ15, §2.3, page 868]. We have an isomorphism of $\mathcal{O}_{U_{i}}$-modules:
$\lambda:\mathscr{H}om_{F_{1}^{*}(S^{<p}T_{U^{\prime}_{i}})}(F_{1}^{*}(S^{<p}T_{U^{\prime}_{i}}),\mathcal{O}_{V_{i}})\xrightarrow{\cong}\mathcal{O}_{V_{i}},\quad\phi\mapsto\phi(1).$
Let $\tilde{\nabla}$ be the connection on $\mathscr{H}om_{F_{1}^{*}(S^{<p}T_{U^{\prime}_{i}})}(F_{1}^{*}(S^{<p}T_{U^{\prime}_{i}}),\mathcal{O}_{V_{i}})$ given by the formula
(~$(\tilde{\nabla}(\partial)(\phi))(\tau)+\phi(\nabla_{T}(\partial)(\tau))=\nabla_{i}(\partial)(\phi(\tau))$
for $\partial\in F_{1}^{*}T_{U^{\prime}_{i}},\phi\in\mathscr{H}om_{F_{1}^{*}(S^{<p}T_{U^{\prime}_{i}})}(F_{1}^{*}(S^{<p}T_{U^{\prime}_{i}}),\mathcal{O}_{V_{i}}),\tau\in F_{1}^{*}(S^{<p}T_{U^{\prime}_{i}}),$ then the $p$-curvature $\psi_{\tilde{\nabla}}$ of $\tilde{\nabla}$ is tautologically zero since
$(\psi_{\tilde{\nabla}}(\partial)(\phi))(\tau)=\psi_{\nabla_{i}}(\partial)(\phi(\tau))-\phi(\psi_{\nabla_{T}}(\partial)(\tau))=0$
hold for all $\partial\in F_{1}^{*}T_{U^{\prime}_{i}},\phi\in\mathscr{H}om_{F_{1}^{*}(S^{<p}T_{U^{\prime}_{i}})}(F_{1}^{*}(S^{<p}T_{U^{\prime}_{i}}),\mathcal{O}_{V_{i}}),\tau\in F_{1}^{*}(S^{<p}T_{U^{\prime}_{i}}).$ This implies that $\nabla_{i}^{\prime}$ has trivial $p$-curvature since ~$\tilde{\nabla}=\nabla_{i}^{\prime}$ via the above isomorphism $\lambda$.
Since the $p$-curvature $\psi_{i}$ is horizontal with respect to $\nabla_{i}$, it is also horizontal with respect to $\nabla^{\prime}_{i}$. Now we want to glue $\{(V_{i},\nabla^{\prime}_{i},\psi_{i})\}_{i\in I}$ to obtain a new flat $G$-bundle $(V^{\prime},\nabla^{\prime})$, of which the $p$-curvature is trivial, together with a $F$-Higgs field $\psi$ horizontal with respect to $\nabla^{\prime}$. We define
@@
which is an automorphism of $V_{ij}$. With the same argument as (3.1), we have